OpenStudy (anonymous):

Find the area of the indicated region: Between y = x^2(x3 + 1)^10 and y = −x(x2 + 1)^10 for x in [0, 1]

OpenStudy (amistre64):

\[\begin{vmatrix}\int_{0}^{1}(f(x)-g(x))\ dx\end{vmatrix}\] where f(x) and g(x) are the given functions

OpenStudy (amistre64):

you would also have to determine if there are any cross overs within the interval to be sure

OpenStudy (amistre64):

x^2(x3 + 1)^10 = −x(x2 + 1)^10 x(x3 + 1)^10 = −(x2 + 1)^10 x= −[(x2 + 1)/(x3 + 1)] ^10 yeah, that does not look fun at all :)

OpenStudy (amistre64):

well, the good news is that they dont intersect between 0 and 1 :) http://www.wolframalpha.com/input/?i=x%5E2%28x3+%2B+1%29%5E10++%3D+%E2%88%92x%28x2+%2B+1%29%5E10